# Enhancing Performance of Explainable AI Models with Constrained Concept Refinement

### v1

@@ Appendix A Related Literature
In the main body of the paper, we provided an overview of related methods in explainable AI. Here, we focus on two approaches that are closely related to and comparable with our proposed method. The first is the CLIP-IP-OMP method introduced by [14], which leverages CLIP embeddings to greedily solve sparse coding problems for each input, subsequently passing the results through a linear layer to generate final predictions. Our method builds upon the same theoretical framework established in [14]. Notably, the baseline model (without concept refinement) considered in Section 3 and Section 4 serves as a differentiable surrogate for CLIP-IP-OMP. The second approach is the label-free CBM (lf-CBM) proposed by [40], which employs the inner product between CLIP embeddings to achieve state-of-the-art results in interpretable models. Specifically, lf-CBM was the first CBM model scaled to ImageNet. A detailed comparison of the performance of our method against these two approaches is provided in Section 4.
@@
Sparse Coding [41] refers to the process of representing a given signal (or vector) as a sparse linear combination of fixed signals, collectively referred to as a dictionary. Dictionary Learning [3] extends Sparse Coding to a bi-variable setting, allowing the dictionary itself to be optimized. The objective is to identify a dictionary capable of generating effective sparse codes for a specific signal distribution. [14] establishes a novel connection between sparse coding and explainable AI, treating concept embeddings as the dictionary. Our proposed method directly optimizes concept embeddings, which draws parallels to dictionary learning, although the two approaches address fundamentally different objectives. Among works on dictionary learning, [36] is particularly relevant, as it optimizes dictionaries for downstream tasks. However, our method distinguishes itself through the constrained concept requirement. Interestingly, [36] aligns with our framework as a differentiable learning module, and integrating this approach with our method presents an intriguing avenue for future work. Another promising direction for future research is leveraging the personalized dictionary learning framework proposed by [33] to extend our approach to Explainable AI in the context of heterogeneous datasets.
Theoretical guarantees. To the best of our knowledge, our work is the first to establish a convergence guarantee in the setting where dictionary atoms are updated for downstream tasks. The most closely related work is [14], which introduced IP-OMP. However, the connection between IP-OMP and column-orthogonal matching pursuit lacks rigor due to the presence of normalization terms (see Theorem 1), making it difficult to prove that IP-OMP minimizes the loss $\mathcal{L}$ or $\mathcal{L}_{m}$ using the optimality conditions of OMP [58]. Another related approach is task-driven dictionary learning [36], which establishes the differentiability of the objective function but does not provide a convergence guarantee.
@@ Appendix B Further Discussion
In this section, we provide further discussion on Theorem 3 and Theorem 4.
Tolerance, Sample Size, and Convergence Speed. Both theorems establish that the population loss converges to zero, provided that the initial dictionary estimate is within $\frac{1}{\sqrt{k}}$ in the column-wise $\|\cdot\|_{1,2}$-norm from the ground-truth dictionary ${\bm{D}}^{*}$. Assuming that a constant fraction of the queries in ${\bm{D}}^{*}$ is utilized to generate each ${\mathbf{x}}^{h}$, i.e., $k=\Omega(n)$, Theorem 4 indicates that each column of ${\bm{D}}^{*}$ can be accurately recovered with a sample size of $m=O(n)$ and a convergence rate of $\alpha=1-\Omega(1)$. Notably, the linear sample size and linear convergence rate is consistent with the best-known results in dictionary learning [5, 34].
Sparsity Level. It is well established in the dictionary learning literature that higher sparsity levels in ground-truth sparse codes (i.e., larger values of $k$ in 3) present greater challenges for recovery; see [5]. Indeed, [5] asserts that sparsity levels beyond $k=\Omega\left(\frac{n}{\log n}\right)$ rarely succeed in practice, even though several approaches have been proposed to handle cases where $k=\Omega\left(n\right)$ [53, 34]. In contrast, our results reveal that the sparsity level $k$ plays a nuanced role when optimizing ${\bm{D}}$ for the downstream task $\mathcal{L}$. On the one hand, a larger $k$ imposes a more stringent initial error bound for both theorems. On the other hand, a larger $k$ in Theorem 4 improves the convergence rate $\tau$ and reduces the required sample size $m$. The former phenomenon is consistent with findings in dictionary learning literature, where exact support recovery becomes more feasible for sparser generative models. A high-level explanation for the latter is that a denser generative model for $\mathbf{\beta}$ provides each ${\mathbf{x}}$ with more information, thereby expediting the training process.
@@ C.1 Proof of Theorem 2
Moreover, if ${\bm{D}}={\bm{D}}^{*}$, we have $S=S^{*}$, where $S^{*}$ is the support of $\beta$ defined in Theorem 2.
We defer the proofs of Lemma 2 and Lemma 3 to Section C.4.2 and Section C.4.3, respectively. Combining the above two lemmas, we obtain:
@@
Given Lemma 3, we have $S=S^{*}$. We can subsequently conclude that:
${\bm{D}}^{*}_{S}{\bm{D}}^{*\top}_{S}{\mathbf{x}}={\bm{D}}^{*}_{S^{*}}{\bm{D}}^{*\top}_{S^{*}}{\mathbf{x}}={\bm{D}}^{*}_{S^{*}}{\bm{D}}^{*\top}_{S^{*}}{\bm{D}}^{*}_{S^{*}}\mathbf{\beta}_{S^{*}}={\bm{D}}^{*}_{S^{*}}\mathbf{\beta}_{S^{*}}={\mathbf{x}},$
@@
The proof of Lemma 4 is presented in Section C.4.4
Our next step is to calculate $\mathcal{L}(f_{\tilde{{\bm{D}}}})$. Note that $\tilde{{\bm{D}}}$ is column-orthogonal according to Lemma 4. Therefore, according to Equation 7, we have $\mathcal{L}(f_{\tilde{{\bm{D}}}})=\|{\bm{D}}^{*}_{S}\tilde{{\bm{D}}}_{S}^{\top}{\mathbf{x}}-{\mathbf{x}}\|^{2}_{2}$. According to Lemma 3, in order to obtain the index set $S$ selected by IP-OMP applied to ${\mathbf{x}}$ and $\tilde{\bm{D}}$, we need to find the indices of the top-$k$ largest values of $|\langle\tilde{{\mathbf{d}}}_{i},{\mathbf{x}}\rangle|$. For $21\leq i\leq k/$ 2, we have:
@@
For $i>k$, we have $|\langle\tilde{{\mathbf{d}}}_{i},{\mathbf{x}}\rangle|=|\langle{\mathbf{d}}^{*}_{i},{\mathbf{x}}\rangle|=0$. As a result, when $\tan\theta<\gamma/\Gamma$, Lemma 3 guarantees that the index set selected by IP-OMP is $S=[k]=S^{*}$. Given this fact, we can calculate $\mathcal{L}(f_{\tilde{{\bm{D}}}})$ as:
@@
This completes the proof of Theorem 2. $\hfill\qed$
C.2 Proof of Theorem 3
@@
Recall that ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$. The equations below rewrite Section 3 for each column $i\in[n]$ of ${\bm{D}}^{(t)}$.
@@
It is easy to verify that the above update rules are equivalent to Section 3. We will use induction to prove this theorem. In particular, define
@@
This two statements combined will establish the correctness of Theorem 3. Indeed, the base case $A(0)$ is trivially satisfied due to Equation 11.
@@
We defer the proof of Lemma 5 to Section C.4.5. To invoke Lemma 5, we have
@@
$=2{\mathbf{x}}\left({\bm{D}}^{*}_{S^{*}}{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-{\bm{D}}^{*}_{S^{*}}\mathbf{\beta}_{S^{*}}\right)^{\top}{\bm{D}}^{*}_{S^{*}}$
$=2{\mathbf{x}}\left({\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\mathbf{\beta}_{S^{*}}\right)^{\top}{\bm{D}}^{*\top}_{S^{*}}{\bm{D}}^{*}_{S^{*}}$
$=2{\mathbf{x}}\left({\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\mathbf{\beta}_{S^{*}}\right)^{\top}.$
@@
$\frac{\partial\mathcal{L}(\tilde{f}_{{\bm{D}}^{(t)}})}{\partial{\mathbf{d}}^{(t)}_{i}}=2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}.$
@@
$\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t+0.5)}_{i}$ $=\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}$
=$\overset{A(t)}{=}\frac{\beta_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(t)}_{i}}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}$
$=\frac{\beta_{i}-{\mathbf{x}}^{\top}\left({\mathbf{d}}^{(t+0.5)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}\right)}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}$
$=\frac{\beta_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(t+0.5)}_{i}}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}-2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}$
@@
In other words, $B(t)$ implies $B(t+0.5)$.
Suppose $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{(0)}_{i}\|_{2}\leq\rho$, then we have ${\mathbf{d}}^{(t+1)}_{i}={\mathbf{d}}^{(t+0.5)}_{i}$, which readily implies $B(t+1)$. On the other hand, if $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{(0)}_{i}\|_{2}>\rho$, we have:
@@
$\|\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t+0.5)}_{i}\|_{2}$ $=\|\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\mathbf{\beta}_{i}){\mathbf{x}}\|_{2}$
@@
We will prove Lemma 6 in Section C.4.6. Now we invoke Lemma 6 with:
@@
$a$ $=\frac{\mathbf{\beta}_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(0)}_{i}}{\|{\mathbf{x}}\|_{2}},$
@@
$a$ $=\frac{\mathbf{\beta}_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(0)}_{i}}{\|{\mathbf{x}}\|_{2}}$ (14)
@@
Now, upon substituting these parameters in Lemma 6, we have $\|{\mathbf{c}}-{\mathbf{a}}\|_{2}\geq\|{\mathbf{b}}-{\mathbf{a}}\|_{2}$, which can be rewritten as:
@@
$=\|{\bm{D}}^{*}_{S^{*}}{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-{\bm{D}}^{*}_{S^{*}}\mathbf{\beta}_{S^{*}}\|^{2}_{2}$
$=\|{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\mathbf{\beta}_{S^{*}}\|^{2}_{2}$
$=\sum_{i\in S^{*}}\left({\mathbf{x}}^{\top}{\mathbf{d}}^{(t)}_{i}-\mathbf{\beta}_{i}\right)^{2}$
@@
which is exactly $B(t+1)$. As a result, we have proved Theorem 3.
C.3 Proof of Theorem 4
@@
We will follow the same index convention defined in Equation 9 and Equation 10, with $\mathcal{L}$ replaced by $\mathcal{L}_{m}$. For each $h\in[m]$, we use $\mathbf{\beta}^{h}$ and $S^{h*}$ to denote the corresponding variables in 3. We define the set $Q_{i}$ as the index set of $h$ such that $i$ is in the support of $\mathbf{\beta}^{h}$:
@@
Given Lemma 5, we have:
$\frac{\partial\mathcal{L}_{m}(\tilde{f}_{{\bm{D}}^{(t)}})}{\partial{\mathbf{d}}^{(t)}_{i}}=\frac{1}{m}\sum_{h\in Q_{i}}2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h},\qquad\forall i\in[n].$
@@
We defer the proof of Lemma 7 to Section C.4.7. Now, we can bound $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}$ as:
$\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}=\left\|{\mathbf{d}}^{(t)}_{i}-\frac{\eta}{m}\sum_{h\in Q_{i}}2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h}-{\mathbf{d}}^{*}_{i}\right\|^{2}_{2}$
$=\|{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}-4\eta\left\langle{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i},\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\rangle+4\eta^{2}\left\|\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\|^{2}_{2}.$ (16)
@@
$\left\langle{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i},\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\rangle$
@@
$\left\|\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\|^{2}_{2}$
@@
$=\frac{1}{m}\sum_{h=1}^{m}\|{\bm{D}}^{(t)\top}_{S^{h*}}{\mathbf{x}}^{h}-\mathbf{\beta}^{h}_{S^{h*}}\|^{2}_{2}$
@@
To complete the proof of Theorem 4, we need to examine the success probability which the above linear convergence occurs with. The only probability statement we use is Lemma 7 and we only invoke it once. So the total success probability is at least:
@@
which is $1-n^{-\omega(1)}$ when $m=\Omega\left(\frac{n^{6}}{\sigma^{2}k^{5}}\right)$. This completes the proof of Theorem 4. $\hfill\qed$

### v2

@@ Appendix A Related Literature
In the main body of the paper, we provided an overview of related methods in explainable AI. Here, we focus on two approaches that are closely related to and comparable with our proposed method. The first is the CLIP-IP-OMP method introduced by (Chattopadhyay et al., 2024), which leverages CLIP embeddings to greedily solve sparse coding problems for each input, subsequently passing the results through a linear layer to generate final predictions. Our method builds upon the same theoretical framework established in (Chattopadhyay et al., 2024). Notably, the baseline model (without concept refinement) considered in Section 3 and Section 4 serves as a differentiable surrogate for CLIP-IP-OMP. The second approach is the label-free CBM (lf-CBM) proposed by (Oikarinen et al., 2023), which employs the inner product between CLIP embeddings to achieve state-of-the-art results in interpretable models. Specifically, lf-CBM was the first CBM model scaled to ImageNet. A detailed comparison of the performance of our method against these two approaches is provided in Section 4.
@@
Sparse Coding (Olshausen & Field, 1997) refers to the process of representing a given signal (or vector) as a sparse linear combination of fixed signals, collectively referred to as a dictionary. Dictionary Learning (Aharon et al., 2006) extends Sparse Coding to a bi-variable setting, allowing the dictionary itself to be optimized. The objective is to identify a dictionary capable of generating effective sparse codes for a specific signal distribution. (Chattopadhyay et al., 2024) establishes a novel connection between sparse coding and explainable AI, treating concept embeddings as the dictionary. Our proposed method directly optimizes concept embeddings, which draws parallels to dictionary learning, although the two approaches address fundamentally different objectives. Among works on dictionary learning, (Mairal et al., 2011) is particularly relevant, as it optimizes dictionaries for downstream tasks. However, our method distinguishes itself through the constrained concept requirement. Interestingly, (Mairal et al., 2011) aligns with our framework as a differentiable learning module, and integrating this approach with our method presents an intriguing avenue for future work. Another promising direction for future research is leveraging the personalized dictionary learning framework proposed by (Liang et al., 2024) to extend our approach to Explainable AI in the context of heterogeneous datasets.
Theoretical guarantees. To the best of our knowledge, our work is the first to establish a convergence guarantee in the setting where dictionary atoms are updated for downstream tasks. The most closely related work is (Chattopadhyay et al., 2024), which introduced IP-OMP. However, the connection between IP-OMP and column-orthogonal matching pursuit lacks rigor due to the presence of normalization terms (see Theorem 2.5), making it difficult to prove that IP-OMP minimizes the loss $\mathcal{L}$ or $\mathcal{L}_{m}$ using the optimality conditions of OMP (Tropp, 2004). Another related approach is task-driven dictionary learning (Mairal et al., 2011), which establishes the differentiability of the objective function but does not provide a convergence guarantee.
@@ Appendix B Further Discussion
In this section, we provide further discussion on Theorem 3.3 and Theorem 3.4.
Tolerance, Sample Size, and Convergence Speed. Both theorems establish that the population loss converges to zero, provided that the initial dictionary estimate is within $\frac{1}{\sqrt{k}}$ in the column-wise $\|\cdot\|_{1,2}$-norm from the ground-truth dictionary ${\bm{D}}^{*}$. Assuming that a constant fraction of the queries in ${\bm{D}}^{*}$ is utilized to generate each ${\mathbf{x}}^{h}$, i.e., $k=\Omega(n)$, Theorem 3.4 indicates that each column of ${\bm{D}}^{*}$ can be accurately recovered with a sample size of $m=O(n)$ and a convergence rate of $\alpha=1-\Omega(1)$. Notably, the linear sample size and linear convergence rate is consistent with the best-known results in dictionary learning (Arora et al., 2015; Liang et al., 2022).
Sparsity Level. It is well established in the dictionary learning literature that higher sparsity levels in ground-truth sparse codes (i.e., larger values of $k$ in 3.2) present greater challenges for recovery; see (Arora et al., 2015). Indeed, (Arora et al., 2015) asserts that sparsity levels beyond $k=\Omega\left(\frac{n}{\log n}\right)$ rarely succeed in practice, even though several approaches have been proposed to handle cases where $k=\Omega\left(n\right)$ (Sun et al., 2016; Liang et al., 2022). In contrast, our results reveal that the sparsity level $k$ plays a nuanced role when optimizing ${\bm{D}}$ for the downstream task $\mathcal{L}$. On the one hand, a larger $k$ imposes a more stringent initial error bound for both theorems. On the other hand, a larger $k$ in Theorem 3.4 improves the convergence rate $\tau$ and reduces the required sample size $m$. The former phenomenon is consistent with findings in dictionary learning literature, where exact support recovery becomes more feasible for sparser generative models. A high-level explanation for the latter is that a denser generative model for $\bm{\beta}$ provides each ${\mathbf{x}}$ with more information, thereby expediting the training process.
@@ Appendix C Proof of Theorems
Moreover, if ${\bm{D}}={\bm{D}}^{*}$, we have $S=S^{*}$, where $S^{*}$ is the support of $\beta$ defined in Theorem 2.6.
We defer the proofs of Lemma C.1 and Lemma C.2 to Section C.4.2 and Section C.4.3, respectively. Combining the above two lemmas, we obtain:
@@
Given Lemma C.2, we have $S=S^{*}$. We can subsequently conclude that:
${\bm{D}}^{*}_{S}{\bm{D}}^{*\top}_{S}{\mathbf{x}}={\bm{D}}^{*}_{S^{*}}{\bm{D}}^{*\top}_{S^{*}}{\mathbf{x}}={\bm{D}}^{*}_{S^{*}}{\bm{D}}^{*\top}_{S^{*}}{\bm{D}}^{*}_{S^{*}}\bm{\beta}_{S^{*}}={\bm{D}}^{*}_{S^{*}}\bm{\beta}_{S^{*}}={\mathbf{x}},$
@@
The proof of Lemma C.3 is presented in Section C.4.4
Our next step is to calculate $\mathcal{L}(f_{\tilde{{\bm{D}}}})$. Note that $\tilde{{\bm{D}}}$ is column-orthogonal according to Lemma C.3. Therefore, according to Equation 8, we have $\mathcal{L}(f_{\tilde{{\bm{D}}}})=\|{\bm{D}}^{*}_{S}\tilde{{\bm{D}}}_{S}^{\top}{\mathbf{x}}-{\mathbf{x}}\|^{2}_{2}$. According to Lemma C.2, in order to obtain the index set $S$ selected by IP-OMP applied to ${\mathbf{x}}$ and $\tilde{\bm{D}}$, we need to find the indices of the top-$k$ largest values of $|\langle\tilde{{\mathbf{d}}}_{i},{\mathbf{x}}\rangle|$. For $21\leq i\leq k/$ 2, we have:
@@
For $i>k$, we have $|\langle\tilde{{\mathbf{d}}}_{i},{\mathbf{x}}\rangle|=|\langle{\mathbf{d}}^{*}_{i},{\mathbf{x}}\rangle|=0$. As a result, when $\tan\theta<\gamma/\Gamma$, Lemma C.2 guarantees that the index set selected by IP-OMP is $S=[k]=S^{*}$. Given this fact, we can calculate $\mathcal{L}(f_{\tilde{{\bm{D}}}})$ as:
@@
This completes the proof of Theorem 2.6. $\hfill\qed$
C.2 Proof of Theorem 3.3
@@
Recall that ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$. The equations below rewrite Equation 5 for each column $i\in[n]$ of ${\bm{D}}^{(t)}$.
@@
It is easy to verify that the above update rules are equivalent to Equation 5. We will use induction to prove this theorem. In particular, define
@@
This two statements combined will establish the correctness of Theorem 3.3 Indeed, the base case $A(0)$ is trivially satisfied due to Equation 12.
@@
We defer the proof of Lemma C.4 to Section C.4.5. To invoke Lemma C.4, we have
@@
$=2{\mathbf{x}}\left({\bm{D}}^{*}_{S^{*}}{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-{\bm{D}}^{*}_{S^{*}}\bm{\beta}_{S^{*}}\right)^{\top}{\bm{D}}^{*}_{S^{*}}$
$=2{\mathbf{x}}\left({\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\bm{\beta}_{S^{*}}\right)^{\top}{\bm{D}}^{*\top}_{S^{*}}{\bm{D}}^{*}_{S^{*}}$
$=2{\mathbf{x}}\left({\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\bm{\beta}_{S^{*}}\right)^{\top}.$
@@
$\frac{\partial\mathcal{L}(\tilde{f}_{{\bm{D}}^{(t)}})}{\partial{\mathbf{d}}^{(t)}_{i}}=2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}.$
@@
$\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t+0.5)}_{i}$ $=\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}$
=$\overset{A(t)}{=}\frac{\beta_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(t)}_{i}}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}$
$=\frac{\beta_{i}-{\mathbf{x}}^{\top}\left({\mathbf{d}}^{(t+0.5)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}\right)}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}$
$=\frac{\beta_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(t+0.5)}_{i}}{\|{\mathbf{x}}\|^{2}_{2}}{\mathbf{x}}-2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}$
@@
In other words, $B(t)$ implies $B(t+0.5)$. Suppose $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{(0)}_{i}\|_{2}\leq\rho$, then we have ${\mathbf{d}}^{(t+1)}_{i}={\mathbf{d}}^{(t+0.5)}_{i}$, which readily implies $B(t+1)$. On the other hand, if $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{(0)}_{i}\|_{2}>\rho$, we have:
@@
$\|\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t+0.5)}_{i}\|_{2}$ $=\|\widehat{{\mathbf{d}}}_{i}-{\mathbf{d}}^{(t)}_{i}+2\eta({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}-\bm{\beta}_{i}){\mathbf{x}}\|_{2}$
@@
We will prove Lemma C.5 in Section C.4.6. Now we invoke Lemma C.5 with:
@@
$a$ $=\frac{\bm{\beta}_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(0)}_{i}}{\|{\mathbf{x}}\|_{2}},$
@@
$a$ $=\frac{\bm{\beta}_{i}-{\mathbf{x}}^{\top}{\mathbf{d}}^{(0)}_{i}}{\|{\mathbf{x}}\|_{2}}$ (15)
@@
Now, upon substituting these parameters in Lemma C.5, we have $\|{\mathbf{c}}-{\mathbf{a}}\|_{2}\geq\|{\mathbf{b}}-{\mathbf{a}}\|_{2}$, which can be rewritten as:
@@
$=\|{\bm{D}}^{*}_{S^{*}}{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-{\bm{D}}^{*}_{S^{*}}\bm{\beta}_{S^{*}}\|^{2}_{2}$
$=\|{\bm{D}}^{(t)\top}_{S^{*}}{\mathbf{x}}-\bm{\beta}_{S^{*}}\|^{2}_{2}$
$=\sum_{i\in S^{*}}\left({\mathbf{x}}^{\top}{\mathbf{d}}^{(t)}_{i}-\bm{\beta}_{i}\right)^{2}$
@@
which is exactly $B(t+1)$. As a result, we have proved Theorem 3.3.
C.3 Proof of Theorem 3.4
@@
We will follow the same index convention defined in Equation 10 and Equation 11, with $\mathcal{L}$ replaced by $\mathcal{L}_{m}$. For each $h\in[m]$, we use $\bm{\beta}^{h}$ and $S^{h*}$ to denote the corresponding variables in 3.2. We define the set $Q_{i}$ as the index set of $h$ such that $i$ is in the support of $\bm{\beta}^{h}$:
@@
Given Lemma C.4, we have:
$\frac{\partial\mathcal{L}_{m}(\tilde{f}_{{\bm{D}}^{(t)}})}{\partial{\mathbf{d}}^{(t)}_{i}}=\frac{1}{m}\sum_{h\in Q_{i}}2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h},\qquad\forall i\in[n].$
@@
We defer the proof of Lemma C.6 to Section C.4.7. Now, we can bound $\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}$ as:
$\|{\mathbf{d}}^{(t+0.5)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}=\left\|{\mathbf{d}}^{(t)}_{i}-\frac{\eta}{m}\sum_{h\in Q_{i}}2({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h}-{\mathbf{d}}^{*}_{i}\right\|^{2}_{2}$
$=\|{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i}\|^{2}_{2}-4\eta\left\langle{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i},\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\rangle+4\eta^{2}\left\|\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\|^{2}_{2}.$ (17)
@@
$\left\langle{\mathbf{d}}^{(t)}_{i}-{\mathbf{d}}^{*}_{i},\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\rangle$
@@
$\left\|\frac{1}{m}\sum_{h\in Q_{i}}({\mathbf{d}}^{(t)\top}_{i}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{i}){\mathbf{x}}^{h}\right\|^{2}_{2}$
@@
$=\frac{1}{m}\sum_{h=1}^{m}\|{\bm{D}}^{(t)\top}_{S^{h*}}{\mathbf{x}}^{h}-\bm{\beta}^{h}_{S^{h*}}\|^{2}_{2}$
@@
To complete the proof of Theorem 3.4, we need to examine the success probability which the above linear convergence occurs with. The only probability statement we use is Lemma C.6 and we only invoke it once. So the total success probability is at least:
@@
which is $1-n^{-\omega(1)}$ when $m=\Omega\left(\frac{n^{6}}{\sigma^{2}k^{5}}\right)$. This completes the proof of Theorem 3.4. $\hfill\qed$